Suppose that the markup of goods prices over marginal cost is 5%, and the wage-setting equation is


W = P(1-u)


where u is the unemployment rate.


The real wage, as determined by the price-setting equation is _______.

The natural rate of unemployment is _______%.
Answer #1

Real Wage = 1/(1+0.05) = 0.9524

Natural Rate of Unemployment = 1- 0.9524 = 0.0476 or 4.76%

answered by: Grace Watson